OpenStudy (anonymous):

Which numbers are greater than 3 × 10−7? Choose all answers that are correct. A. 9 × 10−9 B. 2 × 10−5 C. 6 × 10−7 D. 5 × 10−8 E. 2 × 10−7

OpenStudy (anonymous):

@fufu0331

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

First we need to find their values. Lets start with 3*10-7. 3*10-7 = 30 - 7 = 23

OpenStudy (math&ing001):

So we need to find numbers greater than 23.

OpenStudy (math&ing001):

By the way there aren't any parentheses right ?

OpenStudy (math&ing001):

@deadpool123